# Chemical Signatures of AGB Mass Transfer in Gaia White Dwarf Companions

### v1

@@ Abstract
We present a homogeneous abundance analysis of 160 main-sequence stars in astrometric white-dwarf + main-sequence (WD+MS) binaries with orbits from Gaia DR3. These systems have AU-scale separations and are thought to have undergone mass transfer (MT) when the WD progenitor was an asymptotic giant branch (AGB) star. Using high-resolution spectroscopy, we measure chemical abundances of the MS stars, focusing on s-process elements. Since s-process nucleosynthesis occurs mainly in AGB stars, s-process enhancement in the MS star is a key signature of accretion from an AGB companion. We identify 40 barium dwarfs – 36 of them newly discovered – roughly doubling the known population in astrometric WD+MS binaries and extending it to lower metallicities than previously studied. The s-process abundances show large star-to-star variations that correlate with component masses and with metallicity but not with orbital separation. At the lowest metallicities, three barium dwarfs display strong CH and $\rm C_{2}$ absorption bands, linking them to CEMP-s stars and implying that AGB mass transfer usually leads to strong carbon enhancement at low metallicity. By comparing the observed abundance patterns to AGB nucleosynthesis models, we show that the diversity of s-process enhancements can be explained by variations in donor mass, metallicity, and most importantly, the number of thermal pulses the AGB star experienced before the onset of MT. Variation in the depth of the accretors’ convective envelopes, with which accreted material is diluted, strengthens correlations with MS star mass and metallicity. Our results establish Gaia WD+MS binaries as a powerful laboratory for constraining binary mass-transfer physics and the origins of chemically peculiar stars.
@@ 1 Introduction
The third data release (DR3) of the Gaia mission contained astrometric orbital solutions for over 160,000 binaries (Gaia Collaboration et al., 2023a, b). These solutions were obtained by fitting a Keplerian model to the wobble of unresolved binaries on the plane of the sky due to orbital motion. Long-period orbits produce larger astrometric signals than short-period orbits, but orbits with periods significantly longer than the observational baseline ($\sim 1000\,$d in DR3) cannot be constrained astrometrically. As a result, DR3 astrometric orbits are sensitive to orbital periods ($P_{\rm orb}$) ranging from roughly 100 to 1000 days. Using the DR3 astrometric orbit catalog, S. Shahaf et al. (2024) identified thousands of new WD+MS binary candidates. Most of them host $\sim 0.6\,M_{\odot}$ WDs in AU-scale orbits, suggesting interaction with an asymptotic giant branch (AGB) donor (N. Hallakoun et al., 2024). While this is classically expected to lead to unstable mass transfer, their orbits are significantly wider than previously identified post-common envelope binaries (PCEBs) with $P_{\rm orb}\lesssim 1\,$d (e.g. M. Zorotovic et al., 2010, 2011; S. G. Parsons et al., 2015). Modeling of the selection function reveals that these AU-scale systems are common but were disfavored by most discovery methods before Gaia (N. Yamaguchi et al., 2024b, 2025).
@@
Several recent works have explored possible formation scenarios for AU-scale WD+MS binaries, finding that (1) MT from AGB donors may remain stable over a wider range of initial mass ratios than previously thought (based on “quasi-adiabatic” critical mass ratios derived by e.g. H. Ge et al. 2020; K. D. Temmink et al. 2023 in the general case of evolved donors), and (2) PCEBs may remain wide if MT begins during a thermal pulse of an AGB donor, when its envelope is very loosely bound (e.g. D. Belloni et al. 2024; N. Yamaguchi et al. 2024a, c, who studied the energy budget of the resulting common envelope in the context of the Gaia systems). Still, a formation model that can explain all of the distinctive traits of the Gaia post-MT systems (intermediate orbital separations, non-zero eccentricities, and a narrow WD mass distribution), as well as the sheer abundance of such systems, remains elusive.
@@
One key observational signature of a previous MT phase is the presence of anomalous surface abundances. In particular, barium stars are stars with enhanced abundances of s-process elements, including but not limited to barium. Since the first studied barium stars were giants, the term “barium dwarfs” is commonly used for MS stars enhanced in s-process elements. As s-process nucleosynthesis is expected only to occur during the late stages of stellar evolution on the AGB, the abundances of these pre-AGB stars have been attributed to s-process enhanced material accreted from an AGB companion. Indeed, since their discovery, many barium stars have been confirmed to be in binary systems with WD companions (e.g. R. D. McClure et al., 1980; R. D. McClure, 1983; A. Jorissen & M. Mayor, 1988).
@@
Another closely related population is the carbon enhanced metal-poor (CEMP) stars found in the Galatic halo, with typical abundances of $\mathrm{[Fe/H]}\lesssim-2.0$ and $\mathrm{[C/Fe]}\gtrsim 0.7$ (e.g. S. Rossi et al., 1999; T. C. Beers & N. Christlieb, 2005; W. Aoki et al., 2007a). A large fraction of CEMP stars are also s-process enhanced, with $\mathrm{[Ba/Fe]}\gtrsim 0.5$ (the so-called ‘CEMP-s’ stars; e.g. J. E. Norris et al. 1997; W. Aoki et al. 2007b; K. Jonsell et al. 2006; T. Masseron et al. 2010). Similar to most barium stars, the majority of CEMP-s stars reside in binaries (S. Lucatello et al., 2005) and thus, their s-process enhancement has also been attributed to mass transfer from AGB donors (C. Abate et al., 2015, 2018). However, compared to barium stars, CEMP-s stars are more pristine, meaning it is easier to disentangle signatures of metals originally present in their atmosphere and those of accreted material. They are also older, making them unique probes of binary interactions in the early Universe (e.g. A. Arentsen et al., 2022).
@@
Recently, P. Rekhi et al. (2024, 2025) compiled a sample of 38 systems with s-process enrichment in the Gaia sample of WD+MS binaries using a combination of GALAH archival spectra and their own follow-up spectra. They found that about 30% of WD+MS binaries in the Gaia sample are enhanced in s-process elements, but the s-process abundances are not strongly correlated with the binaries’ orbital parameters. They also identified an anti-correlation between s-process abundances and bulk metallicity in the sample.
@@
In this paper, we expand on these works by targeting a larger sample of WD+MS binaries from the Gaia sample with high-resolution spectroscopy. We target sources with a broad range of metallicities, including several below $\rm[Fe/H]=-2$, and also include several candidate MS + neutron star (NS) binaries. We identify 36 new barium dwarfs, and confirm 4 existing ones in the literature. We find evidence of significant carbon enhancement in three of the most metal-poor barium dwarfs, creating a link to CEMP-s stars. We also present a population model to explain the diversity of s-process abundances found in the Gaia WD+MS binary sample.
@@
The remainder of this paper is organized as follows. In Section 2, we describe target selection for follow-up optical spectroscopy. In Section 3, we describe observations of these targets and the data reduction process. Section 4 details the analysis to constrain atmospheric parameters and measure chemical abundances. In Section 5, we discuss the relationships between the measured abundances and how they vary with the various properties of the systems. We also describe two barium dwarfs orbiting candidate NSs and three with evidence of strong carbon enhancement. In Section 6, we compare our barium dwarfs to several related populations of chemically peculiar stars in the literature. In Section 7, we use theoretical models of s-process yields in the AGB donors and consider the effect of convective mixing in the envelopes of the MS accretors to predict and explain the observed trends and scatter in the barium abundances with the stellar and orbital parameters. Finally, we conclude with a summary of our key findings in Section 8.
2 Target selection
@@
Figure 1: Left: AMRF versus luminous star mass, $M_{1}$, for the full non-class I sample from S. Shahaf et al. (2024) (grey points). Our targets selected for follow-up are shown with colored points, separated by whether they were classified as NCE (pink diamonds) or RCE (blue circles) by S. Shahaf et al. (2024). The dotted and dashed lines are boundaries separating systems in the three classes (Section 2). The shaded regions corresponds to the expected locations of systems hosting dark compact object secondaries with masses between $0.45-0.75\,M_{\odot}$ (yellow, lower) and $1.4-2.1\,M_{\odot}$ (orange, upper), corresponding to typical WD and NS masses. While sources classified as NCE are less likely to be triples, we believe that both the NCE and RCE sources in our follow-up sample are dominated by compact object companions (see text). Right: Color excess against [Fe/H] (X. Zhang et al., 2023) for non-class I systems. The two quantities are anti-correlated, meaning that the NCE sample is biased against metal-poor primaries. We selected several RCE systems lying below the densely populated strip in this space, which are likely to host WDs.
@@
We selected targets from the sample of sources identified by S. Shahaf et al. (2024) as unlikely to be MS+MS binaries. Their selection relies on a quantity called the astrometric mass ratio function (AMRF), which can be defined purely in terms of observables:
@@
where $\alpha$ is the angular photocentric semi-major axis, $\varpi$ is the parallax, $M_{1}$ is the mass of the luminous star, and $P_{\rm orb}$ is the orbital period. As described by S. Shahaf et al. (2019), observed binaries can be grouped into three classes based on their location in the AMRF vs $M_{1}$ plane.
@@
We prioritized NCE candidates in our follow-up. However, many of our targets were selected from a preliminary, unpublished version of the S. Shahaf et al. (2024) sample, and were classified as NCE there but as RCE in the final sample. The preliminary and final samples differed in the adopted metallicities and color excess thresholds. We found that the final NCE selection of S. Shahaf et al. (2024) was conservative: most of the systems classified as RCE class-II in our sample are still likely to host WDs, as evidenced by their mass and eccentricity distributions. Moreover, the fact that many of our class-III targets – which have AMRF values too large to be explained by triples – are nevertheless classified as RCE provides further evidence that many of the objects labeled as RCE are not actually triples.
Simulations from N. Yamaguchi et al. (2025) showed that while the cut on color excess effectively eliminates triples, it also removes about half of all systems hosting WDs in the non-class I sample (see their Figure 5). In particular, they found that the cut is biased against metal-poor systems, likely because the adopted 2 Gyr-old isochrones predict a relation between color and metallicity that is steeper than observed. This means that very few systems with [Fe/H] <0.5<-0.5\,dex enter the NCE sample, suggesting that a metallicity-dependent cut on the color excess may be more appropriate than the flat cut adopted by S. Shahaf et al. (2024).
With this in mind, we selected several metal-poor systems to be observed with HIRES (Section 3) that are in the non-class I sample, but not in the NCE sample of S. Shahaf et al. (2024). As shown in the right panel of Figure 1, we selected metal-poor stars that lie below the average color excess at their metallicties.
@@ 3 Observations
3.1 FEROS
We obtained a total of 219 spectra of 92 targets using the Fiber-fed Extended Range Optical Spectrograph (FEROS; A. Kaufer et al. 1999) instrument installed on the MPG/ESO 2.22.2\,m telescope at the La Silla Observatory (Programs 113.26XB, 114.27SS, and 115.28KE). Exposure times ranged from 600 to 2400 s, depending on the brightness of the target. We used $1\times 1$ binning. This allowed us to achieve a spectral resolution $R\sim 50,000$ over $3860$ to 67706770\,Å. For a single observation, the typical signal-to-noise ratio (SNR) achieved was $\sim 25$ and $15$ above and below 50005000\,Å, respectively. We reduced the raw data using the CERES pipeline (R. Brahm et al., 2017).
3.2 MIKE
We obtained 54 spectra of 24 targets with the Magellan Inamori Kyocera Echelle (MIKE) spectrograph mounted on the Magellan Clay telescope at Las Campanas Observatory (R. Bernstein et al., 2003). We used the $0.7^{\prime\prime}$ slit with $2\times 2$ binning, which resulted in a typical $R\sim 34,000$ and wavelength coverage of 332596803325-9680\,Å. Exposure times ranged from 600 to 2400 s, which yielded typical SNR for a single observation of $\sim 50$ and $40$ above and below 50005000\,Å. We reduced the spectra with the MIKE Pipeline using CarPy (D. D. Kelson et al., 2000; D. D. Kelson, 2003).
3.3 HIRES
We obtained 66 spectra of 61 targets with the High Resolution Echelle Spectrometer (HIRES) on the Keck I telescope. 34 of these targets were observed through the California Planet Search (CPS; A. W. Howard et al. 2010) queue, which allowed for more flexible scheduling. For all observations, we used the C2 decker ($0.86^{\prime\prime}\times 14^{\prime\prime}$ slit) with $3\times 1$ binning, yielding spectra with $R\sim 50,000$. For targets observed through CPS, our exposure times ranged from $120$ to 12001200\,s and we used the HIRESr cross-disperser. This resulted in SNR $\sim 55$ and $35$ above and below 50005000\,Å and wavelength coverage of 364064203640-6420\,Å. For the other targets, we used a fixed exposure time of 1200 s and the HIRESb cross-disperser. This led to SNR $\sim 40$ and $35$ above and below 50005000\,Å, and wavelength coverage 384066403840-6640\,Å.

### v2

@@ Abstract
We present a homogeneous abundance analysis of 160 main-sequence stars in astrometric white-dwarf + main-sequence (WD+MS) binaries with orbits from Gaia DR3. These systems have AU-scale separations and are thought to have undergone mass transfer (MT) when the WD progenitor was an asymptotic giant branch (AGB) star. Using high-resolution spectroscopy, we measure chemical abundances of the MS stars, focusing on s-process elements. Since s-process nucleosynthesis occurs mainly in AGB stars, s-process enhancement in the MS star is a key signature of accretion from an AGB companion. We identify 43 barium dwarfs – 39 of them newly discovered – roughly doubling the known population in astrometric WD+MS binaries and extending it to lower metallicities than previously studied. The s-process abundances show large star-to-star variations that correlate with component masses and with metallicity but not with orbital separation. At the lowest metallicities, three barium dwarfs display strong CH and $\rm C_{2}$ absorption bands, confirming a link between barium stars and CEMP-s stars and implying that AGB mass transfer usually leads to strong carbon enhancement at low metallicity. By comparing the observed abundance patterns to AGB nucleosynthesis models, we show that the diversity of s-process enhancements can be explained by variations in donor mass, metallicity, and most importantly, the number of thermal pulses the AGB star experienced before the onset of MT. Variation in the depth of the accretors’ convective envelopes, with which accreted material is diluted, strengthens correlations with MS star mass and metallicity. Our results establish Gaia WD+MS binaries – which are homogeneously selected and probe shorter orbital periods than previous barium-star samples – as a powerful laboratory for constraining mass transfer physics and the origin of chemically peculiar stars.
@@ I Introduction
The third data release (DR3) of the Gaia mission contained astrometric orbital solutions for over 160,000 binaries (Gaia Collaboration et al., 2023b, a). These solutions were obtained by fitting a Keplerian model to the wobble of unresolved binaries on the plane of the sky due to orbital motion. In the case where the secondary is an optically dark compact object such as a cool WD, the wobble is that of the primary star. Long-period orbits produce larger astrometric signals than short-period orbits, but orbits with periods significantly longer than the observational baseline ($\sim 1000\,$d in DR3) cannot be constrained astrometrically. As a result, DR3 astrometric orbits are sensitive to orbital periods ($P_{\rm orb}$) ranging from roughly 100 to 1000 days. Using the DR3 astrometric orbit catalog, Shahaf et al. (2024) identified thousands of new WD+MS binary candidates. Most of them host $\sim 0.6\,M_{\odot}$ WDs in AU-scale orbits, suggesting interaction with an asymptotic giant branch (AGB) donor (Hallakoun et al., 2024). While this is classically expected to lead to unstable mass transfer, their orbits are significantly wider than previously identified post-common envelope binaries (PCEBs) with $P_{\rm orb}\lesssim 1\,$d (e.g. Zorotovic et al., 2010, 2011; Parsons et al., 2015). Modeling of the selection function reveals that these AU-scale systems are common but were disfavored by most discovery methods before Gaia (Yamaguchi et al., 2024c, 2025).
@@
Several recent works have explored possible formation scenarios for AU-scale WD+MS binaries, finding that (1) MT from AGB donors may remain stable over a wider range of initial mass ratios than previously thought (based on “quasi-adiabatic” critical mass ratios derived by e.g. Ge et al. 2020; Temmink et al. 2023 in the general case of evolved donors), and (2) PCEBs may remain wide if MT begins during a thermal pulse of an AGB donor, when its envelope is very loosely bound (e.g. Belloni et al. 2024; Yamaguchi et al. 2024b, a, who studied the energy budget of the resulting common envelope in the context of the Gaia systems). Still, a formation model that can explain all of the distinctive traits of the Gaia post-MT systems (intermediate orbital separations, non-zero eccentricities, and a narrow WD mass distribution), as well as the sheer abundance of such systems, remains elusive.
@@
One key observational signature of a previous MT phase is the presence of anomalous surface abundances. In particular, barium stars are stars with enhanced abundances of s-process elements, including but not limited to barium. Since the first studied barium stars were giants, the term “barium dwarfs” is commonly used for MS stars enhanced in s-process elements. As s-process nucleosynthesis is expected only to occur during the late stages of stellar evolution on the AGB, the abundances of these pre-AGB stars have been attributed to s-process enhanced material accreted from an AGB companion. Indeed, since their discovery, many barium stars have been confirmed to be in binary systems with WD companions (e.g. McClure et al., 1980; McClure, 1983; Jorissen and Mayor, 1988).
@@
Another closely related population is the carbon enhanced metal-poor (CEMP) stars found in the Galatic halo, with typical abundances of $\mathrm{[Fe/H]}\lesssim-2.0$ and $\mathrm{[C/Fe]}\gtrsim 0.7$ (e.g. Rossi et al., 1999; Beers and Christlieb, 2005; Aoki et al., 2007a). A large fraction of CEMP stars are also s-process enhanced, with $\mathrm{[Ba/Fe]}\gtrsim 0.5$ (the so-called ‘CEMP-s’ stars; e.g. Norris et al. 1997; Aoki et al. 2007b; Jonsell et al. 2006; Masseron et al. 2010). Similar to most barium stars, the majority of CEMP-s stars reside in binaries (Lucatello et al., 2005) and thus, their s-process enhancement has also been attributed to mass transfer from AGB donors (Abate et al., 2015, 2018). However, compared to barium stars, CEMP-s stars are more pristine, meaning it is easier to disentangle signatures of metals originally present in their atmosphere and those of accreted material. They are also older, making them unique probes of binary interactions in the early Universe (e.g. Arentsen et al., 2022). Additionally, dwarf carbon (dC) stars share similar properties to CEMP stars, but as they are less evolved, the origin their carbon enhancement is more narrowly restricted to accretion from a companion (Dearborn et al., 1986; Green et al., 1991; Green, 2013; Downes et al., 2004; Farihi et al., 2018, 2025; Roulston et al., 2019).
@@
Recently, Rekhi et al. (2024, 2025) compiled a sample of 38 systems with s-process enrichment in the Gaia sample of WD+MS binaries using a combination of GALAH archival spectra and their own follow-up spectra. They found that about 30% of WD+MS binaries in the Gaia sample are enhanced in s-process elements, but the s-process abundances are not strongly correlated with the binaries’ orbital parameters. They also identified an anti-correlation between s-process abundances and bulk metallicity in the sample.
@@
In this paper, we expand on these works by targeting a larger sample of WD+MS binaries from the Gaia sample with high-resolution spectroscopy. We target sources with a broad range of metallicities, including several below $\rm[Fe/H]=-2$, and also include several candidate MS + neutron star (NS) binaries. We identify 39 new barium dwarfs, and confirm 4 existing ones in the literature. We find evidence of significant carbon enhancement in three of the most metal-poor barium dwarfs, creating a link to CEMP-s stars. We also present a population model to explain the diversity of s-process abundances found in the Gaia WD+MS binary sample.
A considerable body of previous work has studied the population demographics of chemically peculiar stars (e.g. Escorza et al. 2019; Abate et al. 2015, 2018; Hansen et al. 2016a; Dimoff et al. 2025). These works have confirmed that accretion from an AGB companion can broadly explain the observed abundances and have established evolutionary links between several related classes of chemically peculiar systems. However, previous studies have included only approximate treatments of sample selection functions, making detailed comparisons between models and data challenging. Moreover, many of the previously identified systems have orbital periods longer than 10001000\,d, where wind accretion or wind Roche lobe overflow likely dominate the MT process. Our work thus adds to the existing literature by focusing on systems with orbital solutions from Gaia that have 1001000100-1000\,d periods. This means that they have a homogeneous and well-understood selection function, and their AU-scale orbits suggest that they likely experienced traditional Roche lobe overflow.
@@
The remainder of this paper is organized as follows. In Section II, we describe target selection for follow-up optical spectroscopy. In Section III, we describe observations of these targets and the data reduction process. Section IV details the analysis to constrain atmospheric parameters and measure chemical abundances. In Section V, we discuss the relationships between the measured abundances and how they vary with the various properties of the systems. We also describe two barium dwarfs orbiting candidate NSs and three with evidence of strong carbon enhancement. In Section VI, we compare our barium dwarfs to several related populations of chemically peculiar stars in the literature. In Section VII, we use theoretical models of s-process yields in the AGB donors and consider the effect of convective mixing in the envelopes of the MS accretors to predict and explain the observed trends and scatter in the barium abundances with the stellar and orbital parameters. Finally, we conclude with a summary of our key findings in Section VIII.
II Target selection
@@
Figure 1: Left: AMRF versus luminous star mass, $M_{1}$, for the full non-class I sample from Shahaf et al. (2024) (grey points). Our targets selected for follow-up are shown with colored points, separated by whether they were classified as NCE (pink diamonds) or RCE (blue circles) by Shahaf et al. (2024). The dotted and dashed lines are boundaries separating systems in the three classes (Section II). The shaded regions corresponds to the expected locations of systems hosting dark compact object secondaries with masses between $0.45-0.75\,M_{\odot}$ (yellow, lower) and $1.4-2.1\,M_{\odot}$ (orange, upper), corresponding to typical WD and NS masses. While sources classified as NCE are less likely to be triples, we believe that both the NCE and RCE sources in our follow-up sample are dominated by compact object companions (see text). Right: Color excess against [Fe/H] (Zhang et al., 2023) for non-class I systems. The two quantities are anti-correlated, meaning that the NCE sample is biased against metal-poor primaries. We selected several RCE systems lying below the densely populated strip in this space, which are likely to host WDs.
@@
We selected targets from the sample of sources identified by Shahaf et al. (2024) as unlikely to be MS+MS binaries. Their selection relies on a quantity called the astrometric mass ratio function (AMRF), which can be defined purely in terms of observables:
@@
where $\alpha$ is the angular photocentric semi-major axis, $\varpi$ is the parallax, $M_{1}$ is the mass of the luminous star, and $P_{\rm orb}$ is the orbital period. As described by Shahaf et al. (2019), observed binaries can be grouped into three classes based on their location in the AMRF vs $M_{1}$ plane.
@@
We prioritized NCE candidates in our follow-up. However, many of our targets were selected from a preliminary, unpublished version of the Shahaf et al. (2024) sample, and were classified as NCE there but as RCE in the final sample. The preliminary and final samples differed in the adopted metallicities and color excess thresholds. We found that the final NCE selection of Shahaf et al. (2024) was conservative: most of the systems classified as RCE class-II in our sample are still likely to host WDs, as evidenced by their mass and eccentricity distributions. Moreover, the fact that many of our class-III targets – which have AMRF values too large to be explained by triples – are nevertheless classified as RCE provides further evidence that many of the objects labeled as RCE are not actually triples.
Simulations from Yamaguchi et al. (2025) showed that while the cut on color excess effectively eliminates triples, it also removes about half of all systems hosting WDs in the non-class I sample (see their Figure 5). In particular, they found that the cut is biased against metal-poor systems, likely because the adopted 2 Gyr-old isochrones predict a relation between color and metallicity that is steeper than observed. This means that very few systems with [Fe/H] <0.5<-0.5\,dex enter the NCE sample, suggesting that a metallicity-dependent cut on the color excess may be more appropriate than the flat cut adopted by Shahaf et al. (2024).
As we discuss in Section V.4, the most metal-poor barium dwarfs also tend to be enhanced in carbon. Because carbon forms strong molecular bands in the blue optical (Figure 9), stars that are carbon-rich will appear redder than those that are not (e.g. Ardern-Arentsen et al., 2025). While the magnitude of the color excess resulting from this effect will depend on the exact carbon abundance, this is another caveat of the NCE sample.
With these effects in mind, we selected several metal-poor systems to be observed with HIRES (Section III) that are in the non-class I sample, but not in the NCE sample of Shahaf et al. (2024). As shown in the right panel of Figure 1, we selected metal-poor stars that lie below the average color excess at their metallicties.
@@ III Observations
III.1 FEROS
We obtained a total of 219 spectra of 92 targets using the Fiber-fed Extended Range Optical Spectrograph (FEROS; Kaufer et al. 1999) instrument installed on the MPG/ESO 2.22.2\,m telescope at the La Silla Observatory (Programs 113.26XB, 114.27SS, and 115.28KE). Exposure times ranged from 600 to 2400 s, depending on the brightness of the target. We used $1\times 1$ binning. This allowed us to achieve a spectral resolution $R\sim 50,000$ over $3860$ to 67706770\,Å. For a single observation, the typical signal-to-noise ratio (SNR) achieved was $\sim 25$ and $15$ above and below 50005000\,Å, respectively. We reduced the raw data using the CERES pipeline (Brahm et al., 2017).
III.2 MIKE
We obtained 54 spectra of 24 targets with the Magellan Inamori Kyocera Echelle (MIKE) spectrograph mounted on the Magellan Clay telescope at Las Campanas Observatory (Bernstein et al., 2003). We used the $0.7^{\prime\prime}$ slit with $2\times 2$ binning, which resulted in a typical $R\sim 34,000$ and wavelength coverage of 332596803325-9680\,Å. Exposure times ranged from 600 to 2400 s, which yielded typical SNR for a single observation of $\sim 50$ and $40$ above and below 50005000\,Å. We reduced the spectra with the MIKE Pipeline using CarPy (Kelson et al., 2000; Kelson, 2003).
III.3 HIRES
We obtained 66 spectra of 61 targets with the High Resolution Echelle Spectrometer (HIRES) on the Keck I telescope. 34 of these targets were observed through the California Planet Search (CPS; Howard et al. 2010) queue, which allowed for more flexible scheduling. For all observations, we used the C2 decker ($0.86^{\prime\prime}\times 14^{\prime\prime}$ slit) with $3\times 1$ binning, yielding spectra with $R\sim 50,000$. For targets observed through CPS, our exposure times ranged from $120$ to 12001200\,s and we used the HIRESr cross-disperser. This resulted in SNR $\sim 55$ and $35$ above and below 50005000\,Å and wavelength coverage of 364064203640-6420\,Å. For the other targets, we used a fixed exposure time of 1200 s and the HIRESb cross-disperser. This led to SNR $\sim 40$ and $35$ above and below 50005000\,Å, and wavelength coverage 384066403840-6640\,Å.